OpenStudy (anonymous):

2x2 + 7x - 4x - 6x2. simplify

OpenStudy (anonymous):

just solve first

OpenStudy (anonymous):

here u can see http://www.wolframalpha.com/input/?i=2x2+%2B+7x+-+4x+-+6x2.

OpenStudy (anonymous):

-4x^2 + 3x =0 -4x^2 = -3x

OpenStudy (anonymous):

3x-4x^2

OpenStudy (anonymous):

4x=3 x=3/4

OpenStudy (anonymous):

first if u see that link u will be understanding

OpenStudy (anonymous):

-x2

OpenStudy (anonymous):

comibine alike terms then solve, right? @best.shakir && @Yahoo!

OpenStudy (anonymous):

yes thats wat we did

OpenStudy (anonymous):

okay just making sure. its been a while since ive done this lol